Circular No. 7706

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA SAGITTARII 2001 No. 3 Independent discoveries of an apparent new nova have been received

from A. J. S. Pereira, Carnaxide, Portugal, and from W. Liller, Viña del Mar, Chile. Pereira’s visual discovery was made on Sept. 5.846 UT with 14×100 binoculars, when he estimated the object’s magnitude as 7.0 and noted its strong orange color; Liller found the object at mag 10.0 on a pair of Kodak Technical Pan photographs taken through an orange filter on Sept. 3.979, with a confirming CCD image on Sept. 6.039 showing the star at broadband V = 7.27 (through clouds). Liller provides the following position for the nova: α = 18^{h}11^{m}45^{s}.82 ± 0^{s}.06, δ = −30^{o}30′49′′.9 ± 2′′.1 (equinox 2000.0). Pereira notes that nothing was seen at this position to m_{v} ∼ 8.5 on Sept. 4.8, and to mag 8.0–8.5 on Sept. 1.8 and 2.8. Liller adds that a low-dispersion CCD prism spectrogram obtained of the new nova through clouds shows a strong continuum and only weak Hα emission estimated to be 20 percent brighter than the surrounding continuum. A. Hale, Cloudcroft, NM, reports that a CCD image taken on Sept. 6.1 yields position end figures 46^{s}.01, 50′′.9. Additional visual magnitude estimates: Sept. 5.868, 7.0 (Pereira); 5.880, 6.9 (Pereira); 5.898, 7.0 (Pereira); 5.904, 6.8 (Pereira); 5.928, 6.9 (Pereira); 6.026, 6.9 (D. W. E. Green, Carlisle, MA); 6.13, 7.0 (Hale).

COMET C/2001 A2 (LINEAR) A. Lecacheux, Observatoire de Paris, on behalf of the Odin Team

(http://www.snsb.se/Odin/Odin.html), reports: “The H_{2}O 110-101 line

at 556.936 GHz was observed in C/2001 A2 with high spectral resolution (54 m/s) with the ODIN submillimeter space telescope from June 20 until July 12. From a preliminary analysis, 10′ × 10′ mapping made during July 1.5–2.4 UT shows that the H_{2}O coma is elongated along the comet-sun line; the integrated line intensity inside the 2′ beam peaks at 12 K km s^{−1}, on the main-beam brightness temperature scale. The line width indicates an H_{2}O outflow velocity of 0.7 km/s. The estimated water-production rate for July 1.7–2.0 is 3.8 × 10^{28} molecules/s.”

SUPERNOVA 2001ee IN NGC 2347 On IAUC 7705, lines 3–4, it was meant that M. Armstrong employed

the same program procedures and instrumentation details that T. Boles employed for SN 2001eb.

2001 September 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green